OpenStudy (anonymous):

A particular style of pants costs the retailer $90 per pair. At what price should the retailer mark them so he can sell them at a 10% discount off the selling price ad still make a 30% profit on his cost?

OpenStudy (blacksteel):

Okay, to solve this question we must first determine at what price the retailer has to sell the pants to make a 30% profit. This will be 1.3*90 = $117. Then the pants should sell at this price after a 10% discount is applied to the final price. Thus, we must find the final price x such that 0.9x = 117, or x = 117 / 0.9 = $130.

OpenStudy (anonymous):

130.00 is the answer??

OpenStudy (blacksteel):

Yes.

OpenStudy (anonymous):

thank you

OpenStudy (blacksteel):

No problem!